### v1

# Phase diagram of the interacting Haldane model with spin-dependent sublattice potentials

Abstract
Using the exact-diagonalization (ED) and mean-field (MF) approaches, we study the ground-state phase diagram of the interacting Haldane model on the honeycomb lattice, taking into consideration of the spin-dependent sublattice potentials $\Delta_{\sigma,\alpha}$. Here $\alpha=\text{A}$,B and $\sigma=\uparrow$$\downarrow$ indicate the sublattice and spin components, respectively. If we set $\Delta_{\sigma,\text{A}}=+\Delta$ ($-\Delta$) and $\Delta_{\sigma,\text{B}}$$=-\Delta$ ($+\Delta$) for $\sigma=\uparrow$ ($\downarrow$), the system favors a spin ordered state. On the other hand, introducing the nearest-neighbor Coulomb interaction $V$ can drive the system to be charge ordered. Due to their competition, we find that in both ED and MF approaches, an exotic state with Chern number $C=1$ survives from two locally ordered phases and a topologically ordered phase with $C=2$. In ED method, other properties, including the fidelity metric, the excitation gap and the structure factors, are also used to determine the critical points. In MF method with large enough lattice size, we define the local order parameters and band gap to characterize the phase transitions. The interacting Haldane model and the spin-dependent lattice potential can be realized in an ultracold atom gas, which may serve as a way to detect this intriguing state.
I Introduction
Different from the traditional framework of the Landau-Ginzburg theory that relies on locally defined order parameters resulting from broken symmetries, topological phases have been identified and characterized based on their global, nonlocal propertiesHasan and Kane (2010); Qi and Zhang (2011). Over the past few years, the categorization of topologically ordered states in non-interacting systems has been completed, in terms of different symmetriesChiu et al. (2016); Zhang et al. (2019); Kruthoff et al. (2017); Vergniory et al. (2019); Tang et al. (2019); Schnyder et al. (2008).
On the other hand, interacting topological insulators, which are known as systems encompassing the interplay of topological properties and electronic correlations, have been extensively studied in recent years Rachel (2018); Hohenadler and Assaad (2013). Correlation effects are expected to generate exotic states in the presence of topologically nontrivial conditions, such as the antiferromagnetic topological state in the Bernevig-Hughes-Zhang model with the on-site Hubbard interactionMiyakoshi and Ohta (2013); Yoshida et al. (2013) and the topologically non-trivial phase with $C=1$ ($C$ denotes the Chern number) in the spinful Haldane-Hubbard model on honeycomb latticeHe et al. (2011); Zhu et al. (2014); Wu et al. (2015); Vanhala et al. (2016); Tupitsyn and Prokof’ev (2019); Mertz et al. (2019); Yuan et al. (2023). The origin of this $C=1$ phase is attributed to a spontaneous SU(2) symmetry breaking with one of the spin components in the Hall state and the other one in a localized state. Similar studies include the antiferromagnetic Chern insulator in Kane-Mele-Hubbard model Jiang et al. (2018) and $C=1$ phase in interacting topological models on the square lattice Wang and Qi (2019); Ebrahimkhas et al. (2021). Other attempts like introducing the double exchange processes to the Haldane Hamiltonian can also break the spin symmetry Tran and Tran (2022), while interplays between topology, on-site and nearest-neighbor interactions do not exhibit such phenomenon Wang and Qi (2019); Shao et al. (2021). In experiment, observing these states are still challenging within the field of material science, and thus the trapped cold atoms in controlled optical lattice may provide an alternative and promising way for achieving this purpose. So far, the experimental realization of the topological Haldane model has been reportedJotzu et al. (2014) and the quantum simulations of strongly correlated systems (including the Fermi-Hubbard model) in ultracold Fermi gases can be reviewed in Ref. Hofstetter and Qin (2018); Esslinger (2010); Tarruell and Sanchez-Palencia (2018).
In this paper, inspired by the studies of spin-dependent optical latticeJotzu et al. (2015); Förster et al. (2009); Mandel et al. (2003), we propose an interacting Haldane model on honeycomb lattice with spin-dependent sublattice potentials $\Delta_{\sigma,\alpha}$, where $\sigma=\uparrow,\downarrow$ and $\alpha=$A,B represents the spin and sublattice components, respectively. If one set $\Delta_{\uparrow,\text{A}}=+\Delta$, $\Delta_{\uparrow,\text{B}}$$=-\Delta$, $\Delta_{\downarrow,\text{A}}=-\Delta$ and $\Delta_{\downarrow,\text{B}}$$=+\Delta$, the system favors a spin-density-wave (SDW) state. Meanwhile, we introduce the nearest-neighbor Coulomb interaction $V$ to drive the system into a charge-density-wave (CDW) state. Due to their competitions, an intermediate state can be expected and we thus study its ground-state phase diagram. We find that except for the topologically non-trivial phase with Chern number $C=2$ and two topologically trivial phases (SDW and CDW) with $C=0$, an newly generated topological phase with $C=1$ can be observed in both exact-diagonalization (ED) and mean-field (MF) method. Our finding provide a new version of competitions among topology, electronic correlation and lattice potential to realize the exotic state.
The presentation is organized as follows: we introduce the model, methods and relevant quantities in Sec. II. Sections III presents our results based on the ED and MF approaches. Lastly, a conclusion is given in Sec. IV.
II Model and measurements
The Hamiltonian of the interacting Haldane model with spin-dependent lattice potentials can be writen as
$\hat{H}=\hat{H}_{k}+\hat{H}_{l},$ (1)
where the kinetic part
[…]
and the local part
$\hat{H}_{l}$ $=$ $U\sum_{i}\hat{n}_{i,\uparrow}\hat{n}_{i,\downarrow}+V\sum_{\langle i,j\rangle,\sigma,\sigma^{\prime}}\hat{n}_{i,\sigma}\hat{n}_{i,\sigma^{\prime}}$ (3)
$+$ $\sum_{i,\sigma}\Delta_{\sigma,\alpha}\hat{n}_{i,\sigma}.$
In Eq. (2), $\hat{c}^{\dagger}_{i,\sigma}$ ($\hat{c}_{i,\sigma}$) is the creation (annihilation) operator for an electron at site $i$ with spin $\sigma=\uparrow{\rm or}\downarrow$. $t_{1}$ ($t_{2}$) is the nearest-neighbor (next-nearest-neighbor) hopping constant and the Haldane phase $\phi_{i,j}=\phi$ ($-\phi$) in the clockwise (anticlockwise) loop is introduced to the next-nearest-neighbor hopping terms. In Eq. (3), $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$ is the number operator; $U$ and $V$ are the on-site and nearest-neighbor Coulomb interaction strength, respectively. $\Delta_{\sigma,\alpha}$ is the spin-dependent lattice potential: for spin $\sigma=\uparrow$, $\Delta_{\sigma,A}=+\Delta$ and $\Delta_{\sigma,B}=-\Delta$; for spin $\sigma=\downarrow$, $\Delta_{\sigma,A}=-\Delta$ and $\Delta_{\sigma,B}=+\Delta$.
In what follows, the model in Eq. (1) are named as the extended Haldane-Hubbard model with spin-dependent lattice potentials. Throughout the paper, we set $t_{1}=1$, $t_{2}=0.2$, $\phi=\pi/2$ and focus on the ground-state phase diagram of this model at half-filling.
II.1 Exact diagonalization in real space
The topological invariant is one of the most improtant properties to characterize the topological phase transitions, which can be quantified by the Chern number in our model. Given the twisted boundary conditions (TBCs) Poilblanc (1991), it can be evaluated by Niu et al. (1985),
$C=\int\frac{d\phi_{x}d\phi_{y}}{2\pi{\rm i}}\left(\langle\partial_{\phi_{x}}\Psi|\partial_{\phi_{y}}\Psi\rangle-\langle{\partial_{\phi_{y}}\Psi|\partial_{\phi_{x}}\Psi\rangle}\right),$ (4)
with ${|\Psi\rangle}$ being the ground-state wave function. Here $\phi_{x}$ and $\phi_{y}$ are the twisted phases along two directions. To avoid the integration of the wave function ${|\Psi\rangle}$ with respect to the continuous variables, we instead use a discretized version Fukui et al. (2005); Varney et al. (2011); Zhang et al. (2013) with intervals $\Delta\phi_{x}=2\pi/N_{x}$ and $\Delta\phi_{y}=2\pi/N_{y}$. In what follows, ($N_{x}$, $N_{y}$)=(20, 20) is adopted to calculate the Chern number.
Other properties used to characterize the critical behavior of quantum phase transition include the ground-state fidelity metric $g$, which is defined as Zanardi and Paunković (2006); Campos Venuti and Zanardi (2007); Zanardi et al. (2007)
$g(x,\delta x)\equiv\frac{2}{N}\frac{1-|\langle\Psi(x)|\Psi(x+\delta x)\rangle|}{(\delta x)^{2}},$ (5)
where x represents the parameters V or $\Delta$, and N is the lattice size. $|\Psi(x)\rangle$ [$|\Psi(x+\delta x)\rangle$] is the ground state of $\hat{H}(x)$ [$\hat{H}(x+\delta x)$] and we set $\delta x=10^{-3}$. In addition, the SDW and CDW structure factors can be used to characterize the spin and charge ordered insulators, respectively. Their definitions in a staggered fashion can be written as
$S_{\text{SDW}}=\frac{1}{N}\sum\limits_{i,j}{(-1)}^{\eta}\langle(\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow})(\hat{n}_{j,\uparrow}-\hat{n}_{j,\downarrow})\rangle,$
$S_{\text{CDW}}=\frac{1}{N}\sum\limits_{i,j}{(-1)}^{\eta}\langle(\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow})(\hat{n}_{j,\uparrow}+\hat{n}_{j,\downarrow})\rangle,$ (6)
where $\eta=0$ ($\eta=1$) if sites $i$ and $j$ are in the same (different) sublattice.
II.2 Mean-field method in momentum space
We employ a variational mean-field method in momentum space to contrast the results obtained via the ED method, which has been reported in studying the extended Haldane-Hubbard model without lattice potentials Shao et al. (2021). In our case, by introducing the operators $a_{\textbf{k},\sigma}^{{\dagger}}=\frac{1}{\sqrt{N}}\sum_{i\in\text{A}}c_{i,\sigma}^{{\dagger}}e^{\mathrm{i}\textbf{k}\cdot\textbf{r}_{i}}$ and $b_{\textbf{k},\sigma}^{{\dagger}}=\frac{1}{\sqrt{N}}\sum_{i\in\text{B}}c_{i,\sigma}^{{\dagger}}e^{\mathrm{i}\textbf{k}\cdot\textbf{r}_{i}}$, the Hamiltonian (1) can be expressed as
$\hat{H}=\hat{H}_{\text{0}}+\hat{H}_{\text{I}},$ (7)
where
$\hat{H}_{\text{0}}=\sum_{\textbf{k},\sigma}\left(m_{+,\sigma}(\textbf{k})a_{\textbf{k},\sigma}^{{\dagger}}a_{\textbf{k},\sigma}+m_{-,\sigma}(\textbf{k})b_{\textbf{k},\sigma}^{{\dagger}}b_{\textbf{k},\sigma}\right.$
$\phantom{=\;\;}\left.-t_{1}g(\textbf{k})a_{\textbf{k},\sigma}^{{\dagger}}b_{\textbf{k},\sigma}-t_{1}g^{*}(\textbf{k})b_{\textbf{k},\sigma}^{{\dagger}}a_{\textbf{k},\sigma}\right),$ (8)
and
$\hat{H}_{\text{I}}$ $=\frac{U}{N}\sum_{\textbf{k},\textbf{k'},\textbf{q}}c_{\textbf{k+q},\uparrow}^{{\dagger}}c_{\textbf{k},\uparrow}c_{\textbf{k}^{\prime}-\textbf{q},\downarrow}^{{\dagger}}c_{\textbf{k}^{\prime},\downarrow}$
$+\frac{V}{N}\sum_{\sigma,\sigma^{\prime}}\sum_{\textbf{k},\textbf{k'},\textbf{q}}g(\textbf{q})a_{\textbf{k+q},\sigma}^{{\dagger}}a_{\textbf{k},\sigma}b_{\textbf{k}^{\prime}-\textbf{q},\sigma^{\prime}}^{{\dagger}}b_{\textbf{k}^{\prime},\sigma^{\prime}}.$ (9)
Here $g(\textbf{k})=1+e^{-\mathrm{i}\textbf{k}\cdot\textbf{a}_{1}}+e^{-\mathrm{i}\textbf{k}\cdot\textbf{a}_{2}}$, and we set
$m_{+,\uparrow}(\mathbf{k})$ $=+\Delta+m_{+}(\mathbf{k})$
$m_{+,\downarrow}(\mathbf{k})$ $=-\Delta+m_{+}(\mathbf{k})$
$m_{-,\uparrow}(\mathbf{k})$ $=-\Delta+m_{-}(\mathbf{k})$
$m_{-,\downarrow}(\mathbf{k})$ $=+\Delta+m_{-}(\mathbf{k})$ (10)
with $m_{\pm}(\mathbf{k})=-2t_{2}[\cos({\bf k}\cdot{\bf a}_{1}\mp\phi)+\cos({\bf k}\cdot{\bf a}_{2}\pm\phi)+\cos({\bf k}\cdot({\bf a}_{1}-{\bf a}_{2})\pm\phi)]$.
By decoupling the four-fermion terms in Eq.(9), the mean-field Hamiltonian can be written as
$\hat{H}_{\rm MF}=\hat{H}_{0}+\sum_{\textbf{k}}\psi_{\textbf{k}}^{{\dagger}}\left(\begin{array}[]{cccc}\varepsilon_{\uparrow}^{a}&\xi_{\uparrow\uparrow}(\textbf{k})&\varepsilon_{\uparrow\downarrow}^{a}&\xi_{\uparrow\downarrow}(\textbf{k})\\ \xi^{*}_{\uparrow\uparrow}(\textbf{k})&\varepsilon_{\uparrow}^{b}&\xi^{*}_{\downarrow\uparrow}(\textbf{k})&\varepsilon_{\uparrow\downarrow}^{b}\\ (\varepsilon_{\uparrow\downarrow}^{a})^{*}&\xi_{\downarrow\uparrow}(\textbf{k})&\varepsilon_{\downarrow}^{a}&\xi_{\downarrow\downarrow}(\textbf{k})\\ \xi^{*}_{\uparrow\downarrow}(\textbf{k})&(\varepsilon_{\uparrow\downarrow}^{b})^{*}&\xi^{*}_{\downarrow\downarrow}(\textbf{k})&\varepsilon_{\downarrow}^{b}\end{array}\right)\psi_{\textbf{k}},$
where $\psi^{\dagger}_{\textbf{k}}=[a^{\dagger}_{\textbf{k},\uparrow},b^{\dagger}_{\textbf{k},\uparrow},a^{\dagger}_{\textbf{k},\downarrow},b^{\dagger}_{\textbf{k},\downarrow}]$ is the basis for each lattice momentum ${\bf k}$ and
$\xi_{\sigma\sigma^{\prime}}(\textbf{k})$ $=-\frac{V}{N}\sum_{\textbf{q}}g(\textbf{k}-\textbf{q})\langle b_{\textbf{q},\sigma^{\prime}}^{{\dagger}}a_{\textbf{q},\sigma}\rangle_{\rm MF},$
$\varepsilon_{\sigma}^{a}$ $=Un_{-\sigma}^{a}+3V\sum_{\sigma^{\prime}}n_{\sigma^{\prime}}^{b},$
$\varepsilon_{\sigma}^{b}$ $=Un_{-\sigma}^{b}+3V\sum_{\sigma^{\prime}}n_{\sigma^{\prime}}^{a},$
$\varepsilon_{\uparrow\downarrow}^{a}$ $=-\frac{U}{N}\sum_{\textbf{q}}\langle a_{\textbf{q},\downarrow}^{{\dagger}}a_{\textbf{q},\uparrow}\rangle_{\rm MF},$
$\varepsilon_{\uparrow\downarrow}^{b}$ $=-\frac{U}{N}\sum_{\textbf{q}}\langle b_{\textbf{q},\downarrow}^{{\dagger}}b_{\textbf{q},\uparrow}\rangle_{\rm MF},$ (11)
with densities $n_{\sigma}^{a}=\frac{1}{N}\sum_{\textbf{q}}\langle a_{\textbf{q},\sigma}^{{\dagger}}a_{\textbf{q},\sigma}\rangle_{\rm MF}$ and $n_{\sigma}^{b}=\frac{1}{N}\sum_{\textbf{q}}\langle b_{\textbf{q},\sigma}^{{\dagger}}b_{\textbf{q},\sigma}\rangle_{\rm MF}$. The above mean-field equations can be solved self-consistently by making use of the variational mean-field approach. Once the free energy has converged, the SDW and CDW order parameters can be obtained by
$\mathcal{O}_{\text{SDW}}$ $=$ $\left|\frac{1}{2}\left(\langle\vec{S}_{A}\rangle_{\rm MF}-\langle\vec{S}_{B}\rangle_{\rm MF}\right)\right|,$
$\mathcal{O}_{\text{CDW}}$ $=$ $\left|\left(n^{A}_{\uparrow}+n^{A}_{\downarrow})-(n^{B}_{\uparrow}+n^{B}_{\downarrow}\right)\right|.$ (12)
Here $\vec{S}_{i}=\frac{1}{2}\sum_{\alpha\beta}c_{i\alpha}^{\dagger}\vec{\sigma}_{\alpha\beta}c_{i\beta}$, and $\vec{\sigma}=(\sigma^{x},\sigma^{y},\sigma^{z})$ is the vector of spin-1/2 Pauli matrices. Meanwhile, we use the discrete formulation in its multiband (non-Abelian) version to compute the Chern number Fukui et al. (2005).
III Results and analysis
III.1 Results of the Exact Diagonalization
In our ED calculations, we choose the $12$A cluster whose reciprocal lattice contains the $\Gamma$, K, ${}^{\prime}$ points, and one pair of M points, see more details in Ref. Shao et al. (2021). It has been shown that containing the high-symmetry K and ${}^{\prime}$ points are critical to study the quantum phase transition in interacting Haldane systems Shao et al. (2021); Varney et al. (2010, 2011). Meanwhile, we employ the periodic boundary condition and make use of translational symmetries to reduce the Hilbert space size. Therefore, discussions with regard to the excited states below are restricted to the momentum subspace where the ground state belongs to.
The ($V$, $\Delta$) phase diagrams are shown in Figs. 1(a), 1(b), 1(c) and 1(d) with $U=0.0$, $U=1.0$, $U=2.0$ and $U=3.0$, respectively. The phase transition points (black circles) in Fig. 1 are identified by the positions of the peaks of fidelity metric $g$. Notice that in Fig. 1(a) with $U=0.0$, the results of Chern number matches the phase diagram obtained from the fidelity metric very well, where we use the blue, green and red squares to represent the Chern number $C=2$, $C=1$ and $C=0$, respectively. We can observe from Fig. 1 (a) that the Chern insulator (CI) phase with $C=2$ is in a region with small $V$ and $\Delta$, the spontaneous symmetry breaking SDW or CDW phases with $C=0$ dominates the system when $\Delta$ or $V$ is large enough. Interestingly, an intermediate phase with $C=1$ can be observed surrounded by three expected phases. It is different from the phase diagram in Ref. Vanhala et al. (2016), where the $C=1$ phase is sandwiched between a band insulator and Mott insulator. Notice that their charge ordered phase (band insulator) is governed by the ionic potential $\Delta_{\rm AB}$ which is spin-independent, and their spin ordered phase (Mott insulator) is governed by the on-site Hubbard interaction $U$. While in our case, the charge ordered phase (CDW) is governed by the nearest-neighbor interaction $V$ and the spin ordered phase (SDW) is governed by the spin-dependent lattice potential $\Delta_{\sigma,\alpha}$. We speculate the different competition fashions result in the different locations of $C=1$ phase. In addition, we observe that the $C=1$ phase in Fig. 1(a) can exist with very small $V$, while in small $\Delta$ region ($\Delta\leq 0.5$) it vanishes. This is completely opposite to our results of MF method in Sec. III.2, where the $C=1$ phase can exist in small $\Delta$ region rather than small $V$ region, see Fig. 3(a). We suggest that this is an open issue and expect further studies on it.
Figure 1: Phase diagram in the parametric space ($V$, $\Delta$) of the model (1) based on the results of fidelity metric $g$ with (a) $U=0.0$, (b) $U=1.0$, (c) $U=2.0$ and (d) $U=4.0$. The blue, green and red squares indicate results of Chern number $C=2$, $C=1$ and $C=0$, respectively. The black dashed lines in (a) and (d) denote the parameters we choose to show more details below.
When $U=1.0$ in Fig. 1(b), however, results of Chern number do not match the phase diagram obtained from the fidelity metric. Similar discrepancies between Chern number and fidelity metric have also been discussed in Ref. Yuan et al. (2023) and they found that increasing $(N_{x},N_{y})$ from (6, 6) to (20, 20) can effectively reduce such discrepancies. Remember that $N_{x}$ and $N_{y}$ come from the discretized method to calculate Chern number, see the relevant discussions in Sec. II. While here we already use $(N_{x}=20,N_{y}=20)$ to calculate the Chern number, and results of $C=1$ (green points) can not even be observed in Figs. 1(c) and 1(d) with $U=2.0$ and $U=3.0$, respectively. We further use $(N_{x}=60,N_{y}=60)$ to check several points in Fig. 1(d) but with no finding of $C=1$. On the other hand, the area of the intermediate phase identified by the fidelity metric decreases but does not vanishes as $U$ increases from $0.0$ to $3.0$. (We must have a statement here that in our ED discussions, the $C=1$ phase is equivalent to the intermediate phase when $U=0.0$, and they are not equivalent when $U$ is finite). The MF results in Sec. III.2 also show the existence of $C=1$ in the case of $U=3.0$. We thus suggest the issue of whether the $C=1$ phase can exist with $U$ increased to 3.0 may still be open. This issue is worth of further study by other methods like Monte Carlo and density matrix renormalization group because in materials $U$ is generally larger than $V$. Back to Fig. 1(a), the situation with $U=0.0$ and finite $V$ may still be realized in ultracold atom systems due to their advances in parameter simulations.
Figure 2: (a)(e) Four lowest-lying energy levels $E_{\alpha}$, (b)(f) the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$, (c)(g) the structure factors $S_{\text{SDW/CDW}}$, and (d)(h) the fidelity metric $g$ of the model (1) with $V=0.4$ on the left panels and $\Delta=1.4$ on the right panels. The on-site interaction $U=0$ and the parameters are corresponding to the balck dashed lines in Fig. 1(a).
To detail the critical behaviors of the phase diagram, we select two balck dashed lines with $V=0.4$ and $\Delta=1.4$ in Fig. 1(a) to calculate the relevant properties. The first four lowest-lying energy levels $E_{\alpha}$ ($\alpha=0,1,2,3$ and $\alpha=0$ represents the ground state) are obtained by employing the Arnoldi Lehoucq et al. (1997) method, see Figs. 2(a) and 2(e). In Fig. 2(a) with $V=0.4$, the system crosses the $C=2$, $C=1$ and SDW phases as $\Delta$ increases from $0.6$ to $1.4$, and two level crossings between the ground state and one excited state can be observed at $\Delta\approx 0.9$ and $\Delta\approx 1.2$. In Fig. 2(e) with $\Delta=1.4$, the system crosses the SDW, $C=1$ and CDW phases as $V$ increases from $0.0$ to $2.0$, and two level crossings locates at $V\approx 0.75$ and $V\approx 1.25$. It may be more intuitive to exhibit the excitation gaps $\Delta_{\rm ex}^{1}=E_{1}-E_{0}$ and $\Delta_{\rm ex}^{2}=E_{2}-E_{0}$, see Figs. 2(b) and 2(f). We can find that the excitation gaps exhibit local minimum values at the critical points where level crossings occur. It should be pointed out that in Fig. 2(f), $\Delta_{\rm ex}^{1}$ becomes very small when $V>1.25$, which is due to the almost degenerate ground state and first excited state in CDW phase. Another feature to characterize the phase transitions is the change of the CDW or SDW structure factor ($S_{\text{CDW}}$ or $S_{\text{SDW}}$) in Figs. 2(c) and 2(g). Values of $S_{\text{CDW}}$ ($S_{\text{SDW}}$) in $C=1$ phase is larger than those in $C=2$ phase but smaller than those in CDW (SDW) phase, which indicating that $C=1$ is an intermediate state as a result of the interplay between topology, electronic correlations and lattice potentials. In addition, sharp peaks of the ground-state fidelity metrics $g$ in Figs. 2(d) and 2(h) can be also observed and used to characterize the critical points. Finally, notice that the left and right panels in Fig. 2 are corresponding to the results of $V=0.4$ and $\Delta=1.4$, respectively.
For $U=3.0$ in Fig. 1(d), similar calculations and discussions along the black dashed lines of $\Delta=0.4$ and $\Delta=0.6$ are detailed in Appendix A.
III.2 Results of the Mean-Field Method
To contrast the exact results on small lattice above, we now report the outcomes of the MF method. By implementing a variational mean-field approximation, the lattice size can be enlarged significantly. We use the $180\times 180$ lattice to calculate the band gap and the order parameters ($\mathcal{O}_{\text{CDW}}$ and $\mathcal{O}_{\text{SDW}}$), while a $30\times 30$ lattice is adopted for the calculations of Chern number. Even though their lattice size are different, critical points identified by the Chern number and other properties are consistent, as will be discussed in Fig. 4. Phase diagrams with regard to the parameters $V$ and $\Delta$ are shown in Figs. 3(a), 3(b), 3(c) and 3(d) with $U=0.0$, $U=1.0$, $U=2.0$ and $U=3.0$, respectively. Compared to the phase diagrams of ED in Fig. 1, similar features can be observed: SDW and CDW dominate the system for large $\Delta$ and $V$, respectively, leaving the $C=2$ phase in small ($V$, $\Delta$) region; the $C=1$ phase are surrounded by the above three phases. Differently, the increasing interaction $U$ suppresses but does not eliminate the $C=1$ state.
Figure 3: Mean-field phase diagram in the parametric space ($V$, $\Delta$) of the model (1) based on the results of Chern number $C$ and the order parameters ($\mathcal{O}_{\text{CDW}}$ and $\mathcal{O}_{\text{SDW}}$), with (a) $U=0.0$, (b) $U=1.0$, (c) $U=2.0$ and (d) $U=4.0$.
Taking the $U=0.0$ case in Fig. 3(a) as an example, we show in Figs. 4(a) and 4(b) the CDW and SDW order parameters ($\mathcal{O}_{\text{CDW}}$ and $\mathcal{O}_{\text{SDW}}$), respectively, as a function of $V$ and $\Delta$. We find that $\mathcal{O}_{\text{CDW}}$ vanishes in SDW and $C=2$ phases, while the $C=1$ phase can be regarded as a zone of transition for $\mathcal{O}_{\text{CDW}}$ decreases from $2$ to 00. Notice that the phase boundary between SDW and CDW can only be identified from the order parameters, and other phase boundaries are characterized by the results of Chern number, see Fig. 4(d). Even though there are some defect points in Fig. 4(d) which may be due to the severe quantum fluctuation in the vicinity of the critical points, the phase boundaries can be readily identified. For $\mathcal{O}_{\text{SDW}}$ in Fig. 4(b), the $C=1$ phase is still a transition region with finite values but smaller that those in SDW phase. These features are similar to the structure factors of ED results shown in Figs. 2(c) and 2(g). To contrast the excitation gaps in ED results, we show the band gap defined as $\Delta(\textbf{k})$=min [$E_{2}(\textbf{k}^{\prime})-E_{1}(\textbf{k}^{\prime})$] in Fig. 4(c). The rapid drop or closing of the gap size can be observed at the critical values except for the phase boundary between SDW and CDW, where only discontinuity of gap size appears instead. This can be connected with the general picture that the change of a topological invariant is always accompanied by a single-particle gap closing. For the phase transition from SDW to CDW side, the Chern number does not change ($C=0$). That may be a reason to explain the discontinuity of gap size instead of gap closing, and remember that they are all gapped phases with different gap sizes.
Figure 4: Contour plots of (a) the CDW order parameter, (b) the SDW order parameter, (c) the band gap and (d) the Chern number, as a function of $V$ and $\Delta$. The mean-field approach is adopted for the calculations of the model (1) with $U=0.0$.
IV Summary and Discussion
To summarize, we studied the interacting spinful Haldane model at half-filling on the honeycomb lattice with spin-dependent sublattice potentials. By employing the exact-diagonalization (ED) and mean-field (MF) method, we obtained the similar ground-state phase diagrams with regard to the nearest-neighbor interaction $V\in[0,3]$ and sublattice potential difference $\Delta\in[0,3]$: the CDW and SDW phases are in the large $V$ and $\Delta$ region, respectively; the $C=2$ phase dominates the system when both $V$ and $\Delta$ are small enough; a intermediate phase with $C=1$ are surrounded by the above three phases. Except for the change of topological invariant, i.e., the Chern number, other features like the closure of excitation gap, change of structure factors and peak of the fidelity metric in ED results are also observed. As a contrast, the mean-field Chern number, band gap and local order parameters are also studied, which makes the existence of such $C=1$ phase more evident.
On the other hand, the interaction $U$ suppresses the $C=1$ phase but leaving an open question of whether it can exist with $U$ increased to $3.0$. In ED, $C=1$ can not be found when $U=3.0$ even though an intermediate phase can still be identified by results of the fidelity metric. In MF, $C=1$ phase preserves for $U=3.0$ even though its area is very small. Further studies on this interesting issue are thus expected.
Acknowledgements.
C. S. acknowledges support from the National Natural Science Foundation of China (NSFC; Grant No. 12104229) and the Fundamental Research Funds for the Central Universities (Grant No. 30922010803). H.-G. L. acknowledges support from NSFC (Grants No. 11834005 and No. 12247101), and the National Key Research and Development Program of China (Grant No. 2022YFA1402704).
Appendix A details of some properties with $U=3$
Figure 5: (a)(e) Four lowest-lying energy levels $E_{\alpha}$, (b)(f) the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$, (c)(g) the structure factors $S_{\text{SDW/CDW}}$, and (d)(h) the fidelity metric $g$ of the model (1) with $\Delta=0.4$ on the left panels and $\Delta=0.6$ on the right panels. The on-site interaction $U=3$ and the parameters are corresponding to the balck dashed lines in Fig. 1(d).
In the case of $U=3.0$, we select $\Delta=0.4$ (left panel) and $\Delta=0.6$ (right panel) to present the relevant properties in Fig. 5. Remember that we do not observe the results of $C=1$ in Fig. 1(d) with $U=3.0$. The line of $\Delta=0.4$ ($\Delta=0.6$) crosses the $C=2$ phase (SDW phase), the intermediate phase and the CDW phase as $V$ increases from $1.0$ to $2.0$ [see black dashed lines in Fig. 1(d)]. From the ground-state fidelity metrics $g$ in Figs. 5(d) and 5(h), we can find the sharp peaks between the intermediate and the CDW phases. Accompanied with the shape peaks, level crossings between the ground state and second excited state [Figs. 5(a) and 5(e)], minimum of the second excitation gap $\Delta_{\text{ex}}^{(2)}$ [Figs. 5(b) and 5(f)] and sudden changes of structure factors [Figs. 5(c) and 5(g)] can be observed. Notice that in CDW phase the ground state and first excited state are almost degenerate. On the other hand, only “hump”s can be observed at the phase boundaries between the $C=2$ (or SDW) and the intermediate phase. Near the positions of the “hump”s, evolutions of the low-energy levels are more complicated and changes of structure factors are more smooth. Such features have also been discussed in Refs. Shao et al. (2021); Yuan et al. (2023), where its reason is attributed to the finite-size effect and it can be eliminated by using different clusters or twisted boundary conditions.

### v3

@@
 Abstract
-Using the exact-diagonalization (ED) and mean-field (MF) approaches, we study the ground-state phase diagram of the interacting Haldane model on the honeycomb lattice, taking into consideration of the spin-dependent sublattice potentials $\Delta_{\sigma,\alpha}$.
-Here $\alpha=\text{A}$,B and $\sigma=\uparrow$$\downarrow$ indicate the sublattice and spin components, respectively.
-If we set $\Delta_{\sigma,\text{A}}=+\Delta$ ($-\Delta$) and $\Delta_{\sigma,\text{B}}$$=-\Delta$ ($+\Delta$) for $\sigma=\uparrow$ ($\downarrow$), the system favors a spin ordered state.
-On the other hand, introducing the nearest-neighbor Coulomb interaction $V$ can drive the system to be charge ordered.
-Due to their competition, we find that in both ED and MF approaches, an exotic state with Chern number $C=1$ survives from two locally ordered phases and a topologically ordered phase with $C=2$.
-In ED method, other properties, including the fidelity metric, the excitation gap and the structure factors, are also used to determine the critical points.
-In MF method with large enough lattice size, we define the local order parameters and band gap to characterize the phase transitions.
-The interacting Haldane model and the spin-dependent lattice potential can be realized in an ultracold atom gas, which may serve as a way to detect this intriguing state.
+Using the exact-diagonalization (ED) and mean-field (MF) approaches, we investigate the ground-state phase diagram of the interacting Haldane model on the honeycomb lattice, incorporating spin-dependent sublattice potentials $\Delta_{\sigma,\alpha}$.
+Here $\alpha=\text{A}$,B and $\sigma=\uparrow$$\downarrow$ denote the sublattice and spin components, respectively.
+Setting $\Delta_{\sigma,\text{A}}=+\Delta$ ($-\Delta$) and $\Delta_{\sigma,\text{B}}$$=-\Delta$ ($+\Delta$) for $\sigma=\uparrow$ ($\downarrow$) results in the system favoring a spin ordered state.
+Conversely, introducing the nearest-neighbor Coulomb interaction can induce charge ordering in the system.
+Due to the competition between these factors, we observe that in both ED and MF approaches, an exotic state with Chern number $C=1$ survives amidst two locally ordered phases and a topologically ordered phase with $C=2$.
+In the ED method, various properties, such as the fidelity metric, the excitation gaps and the structure factors, are employed to identify critical points.
+In the MF method, using a sufficiently large lattice size, we define the local order parameters and band gaps to characterize the phase transitions.
+The interacting Haldane model and the spin-dependent lattice potential may be experimentally realized in an ultracold atom gas, providing a potential means to detect this intriguing state.
 I Introduction
-Different from the traditional framework of the Landau-Ginzburg theory that relies on locally defined order parameters resulting from broken symmetries, topological phases have been identified and characterized based on their global, nonlocal propertiesHasan and Kane (2010); Qi and Zhang (2011).
-Over the past few years, the categorization of topologically ordered states in non-interacting systems has been completed, in terms of different symmetriesChiu et al.
+In contrast to the traditional framework of the Landau-Ginzburg theory, which relies on locally defined order parameters resulting from broken symmetries, topological phases have been identified and characterized based on their global, nonlocal propertiesHasan and Kane (2010); Qi and Zhang (2011).
+Over the past few years, the categorization of topologically ordered states in non-interacting systems has been completed, considering various symmetriesChiu et al.
 (2016); Zhang et al.
@@
 (2008).
-On the other hand, interacting topological insulators, which are known as systems encompassing the interplay of topological properties and electronic correlations, have been extensively studied in recent years Rachel (2018); Hohenadler and Assaad (2013).
-Correlation effects are expected to generate exotic states in the presence of topologically nontrivial conditions, such as the antiferromagnetic topological state in the Bernevig-Hughes-Zhang model with the on-site Hubbard interactionMiyakoshi and Ohta (2013); Yoshida et al.
-(2013) and the topologically non-trivial phase with $C=1$ ($C$ denotes the Chern number) in the spinful Haldane-Hubbard model on honeycomb latticeHe et al.
+On the other hand, there has been extensive research on the interacting topological insulators, which involve the interplay between topological properties and electronic correlations Rachel (2018); Hohenadler and Assaad (2013).
+Correlation effects are expected to give rise to exotic states in the presence of topologically nontrivial conditions.
+Examples include the antiferromagnetic topological state in the Bernevig-Hughes-Zhang model with the on-site Hubbard interactionMiyakoshi and Ohta (2013); Yoshida et al.
+(2013) and the topologically non-trivial phase with $C=1$ ($C$ denoting the Chern number) in the spinful Haldane-Hubbard model on honeycomb latticeHe et al.
 (2011); Zhu et al.
@@
 (2019); Yuan et al.
+(2023); He et al.
 (2024).
-The origin of this $C=1$ phase is attributed to a spontaneous SU(2) symmetry breaking with one of the spin components in the Hall state and the other one in a localized state.
-Similar studies include the antiferromagnetic Chern insulator in Kane-Mele-Hubbard model Jiang et al.
+The origin of this $C=1$ phase is attributed to a spontaneous SU(2) symmetry breaking, with one spin component in the Hall state and the other in a localized state.
+Similar investigations include the antiferromagnetic Chern insulator in Kane-Mele-Hubbard model Jiang et al.
 (2018) and $C=1$ phase in interacting topological models on the square lattice Wang and Qi (2019); Ebrahimkhas et al.
 (2021).
-Other attempts like introducing the double exchange processes to the Haldane Hamiltonian can also break the spin symmetry Tran and Tran (2022), while interplays between topology, on-site and nearest-neighbor interactions do not exhibit such phenomenon Wang and Qi (2019); Shao et al.
+Additionally, introducing disorder to the interacting Haldane model Silva et al.
+or double exchange processes to the Haldane Hamiltonian Tran and Tran (2022) can also break the spin symmetry.
+However, interplays between topology, on-site, and nearest-neighbor interactions do not exhibit such phenomenon Wang and Qi (2019); Shao et al.
 (2021).
-In experiment, observing these states are still challenging within the field of material science, and thus the trapped cold atoms in controlled optical lattice may provide an alternative and promising way for achieving this purpose.
-So far, the experimental realization of the topological Haldane model has been reportedJotzu et al.
-(2014) and the quantum simulations of strongly correlated systems (including the Fermi-Hubbard model) in ultracold Fermi gases can be reviewed in Ref.
+In experimental settings, observing these states remains challenging in realistic materials.
+Trapped cold atoms may offer an alternative and promising approach to achieving this purpose.
+Notably, the experimental realization of the topological Haldane model has been reportedJotzu et al.
+(2014).
+Quantum simulations of strongly correlated systems in ultracold Fermi gases, including the Fermi-Hubbard model, have also been reviewed in Ref.
 Hofstetter and Qin (2018); Esslinger (2010); Tarruell and Sanchez-Palencia (2018).
-In this paper, inspired by the studies of spin-dependent optical latticeJotzu et al.
+In this paper, drawing inspiration from studies on spin-dependent optical latticeJotzu et al.
 (2015); Förster et al.
 (2009); Mandel et al.
-(2003), we propose an interacting Haldane model on honeycomb lattice with spin-dependent sublattice potentials $\Delta_{\sigma,\alpha}$, where $\sigma=\uparrow,\downarrow$ and $\alpha=$A,B represents the spin and sublattice components, respectively.
-If one set $\Delta_{\uparrow,\text{A}}=+\Delta$, $\Delta_{\uparrow,\text{B}}$$=-\Delta$, $\Delta_{\downarrow,\text{A}}=-\Delta$ and $\Delta_{\downarrow,\text{B}}$$=+\Delta$, the system favors a spin-density-wave (SDW) state.
-Meanwhile, we introduce the nearest-neighbor Coulomb interaction $V$ to drive the system into a charge-density-wave (CDW) state.
-Due to their competitions, an intermediate state can be expected and we thus study its ground-state phase diagram.
-We find that except for the topologically non-trivial phase with Chern number $C=2$ and two topologically trivial phases (SDW and CDW) with $C=0$, an newly generated topological phase with $C=1$ can be observed in both exact-diagonalization (ED) and mean-field (MF) method.
-Our finding provide a new version of competitions among topology, electronic correlation and lattice potential to realize the exotic state.
-The presentation is organized as follows: we introduce the model, methods and relevant quantities in Sec.
+(2003), we propose an interacting Haldane model on honeycomb lattice with spin-dependent sublattice potentials $\Delta_{\sigma,\alpha}$, where $\sigma=\uparrow,\downarrow$ and $\alpha=$A,B represent the spin and sublattice indices, respectively.
+Setting $\Delta_{\sigma,\text{A}}=+\Delta$ ($-\Delta$) and $\Delta_{\sigma,\text{B}}$$=-\Delta$ ($+\Delta$) for $\sigma=\uparrow$ ($\downarrow$) leads to the system favoring a staggered spin order (SSO) state.
+While introducing the nearest-neighbor Coulomb interaction $V$ drives the system into a staggered charge order (SCO) state.
+Due to their competitions, an intermediate state is expected and we explore the ground-state phase diagram of this model.
+Our findings reveal that, in addition to the topologically nontrivial phase with Chern number $C=2$ and two topologically trivial phases (SSO and SCO) with $C=0$, an newly generated phase with $C=1$ can be observed in both exact-diagonalization (ED) and mean-field (MF) methods.
+Various properties, including the excitation gaps, the structure factors, and the fidelity metric in ED, as well as the band gap and local order parameters in MF, are obtained to characterize the phase transitions.
+Especially, an effective potential differences for the two spin species in MF method are defined to analyze the origin of the $C=1$ phase.
+Our finding presents a perspective on the interplay between topology, electronic correlation and lattice potential, shedding light on the realization of exotic states.
+The presentation is structured as follows: we introduce the model, methods and relevant quantities in Sec.
 II.
 Sections III presents our results based on the ED and MF approaches.
-Lastly, a conclusion is given in Sec.
+Lastly, a conclusion is provided in Sec.
 IV.
@@
 and the local part
-$\hat{H}_{l}$ $=$ $U\sum_{i}\hat{n}_{i,\uparrow}\hat{n}_{i,\downarrow}+V\sum_{\langle i,j\rangle,\sigma,\sigma^{\prime}}\hat{n}_{i,\sigma}\hat{n}_{i,\sigma^{\prime}}$ (3)
-$+$ $\sum_{i,\sigma}\Delta_{\sigma,\alpha}\hat{n}_{i,\sigma}.$
+$\hat{H}_{l}$ $=$ $U\sum_{i}(\hat{n}_{i,\uparrow}-\frac{1}{2})(\hat{n}_{i,\downarrow}-\frac{1}{2})$ (3)
+$+$ $V\sum_{\langle i,j\rangle}(\hat{n}_{i}-1)(\hat{n}_{j}-1)+\sum_{i,\sigma}\Delta_{\sigma,\alpha}\hat{n}_{i,\sigma}.$
 In Eq.
-(2), $\hat{c}^{\dagger}_{i,\sigma}$ ($\hat{c}_{i,\sigma}$) is the creation (annihilation) operator for an electron at site $i$ with spin $\sigma=\uparrow{\rm or}\downarrow$.
+(2), $\hat{c}^{\dagger}_{i,\sigma}$ ($\hat{c}^{\phantom{}}_{i,\sigma}$) is the creation (annihilation) operator for an electron at site $i$ with spin $\sigma=\uparrow{\rm or}\downarrow$.
 $t_{1}$ ($t_{2}$) is the nearest-neighbor (next-nearest-neighbor) hopping constant and the Haldane phase $\phi_{i,j}=\phi$ ($-\phi$) in the clockwise (anticlockwise) loop is introduced to the next-nearest-neighbor hopping terms.
 In Eq.
-(3), $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$ is the number operator; $U$ and $V$ are the on-site and nearest-neighbor Coulomb interaction strength, respectively.
+(3), $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}^{\phantom{}}$ and $\hat{n}_{i}=\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow}$; $U$ and $V$ are the on-site and nearest-neighbor Coulomb interactions, respectively.
 $\Delta_{\sigma,\alpha}$ is the spin-dependent lattice potential: for spin $\sigma=\uparrow$, $\Delta_{\sigma,A}=+\Delta$ and $\Delta_{\sigma,B}=-\Delta$; for spin $\sigma=\downarrow$, $\Delta_{\sigma,A}=-\Delta$ and $\Delta_{\sigma,B}=+\Delta$.
 In what follows, the model in Eq.
-(1) are named as the extended Haldane-Hubbard model with spin-dependent lattice potentials.
+(1) is named as the extended Haldane-Hubbard model with spin-dependent lattice potentials.
 Throughout the paper, we set $t_{1}=1$, $t_{2}=0.2$, $\phi=\pi/2$ and focus on the ground-state phase diagram of this model at half-filling.
@@
 In what follows, ($N_{x}$, $N_{y}$)=(20, 20) is adopted to calculate the Chern number.
-Other properties used to characterize the critical behavior of quantum phase transition include the ground-state fidelity metric $g$, which is defined as Zanardi and Paunković (2006); Campos Venuti and Zanardi (2007); Zanardi et al.
+Other properties used to characterize the critical behavior include the ground-state fidelity metric $g$, which is defined as Zanardi and Paunković (2006); Campos Venuti and Zanardi (2007); Zanardi et al.
 (2007)
@@
 $|\Psi(x)\rangle$ [$|\Psi(x+\delta x)\rangle$] is the ground state of $\hat{H}(x)$ [$\hat{H}(x+\delta x)$] and we set $\delta x=10^{-3}$.
-In addition, the SDW and CDW structure factors can be used to characterize the spin and charge ordered insulators, respectively.
-Their definitions in a staggered fashion can be written as
-$S_{\text{SDW}}=\frac{1}{N}\sum\limits_{i,j}{(-1)}^{\eta}\langle(\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow})(\hat{n}_{j,\uparrow}-\hat{n}_{j,\downarrow})\rangle,$
-$S_{\text{CDW}}=\frac{1}{N}\sum\limits_{i,j}{(-1)}^{\eta}\langle(\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow})(\hat{n}_{j,\uparrow}+\hat{n}_{j,\downarrow})\rangle,$ (6)
-where $\eta=0$ ($\eta=1$) if sites $i$ and $j$ are in the same (different) sublattice.
+In addition, the SSO and SCO structure factors can be used to characterize the spin and charge ordered states, respectively, and their definitions in a staggered fashion can be written as
+$S_{\rm SSO}=\frac{1}{N}\sum\limits_{i,j}{(-1)}^{\eta}\langle(\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow})(\hat{n}_{j,\uparrow}-\hat{n}_{j,\downarrow})\rangle,$
+$S_{\rm SCO}=\frac{1}{N}\sum\limits_{i,j}{(-1)}^{\eta}\langle(\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow})(\hat{n}_{j,\uparrow}+\hat{n}_{j,\downarrow})\rangle,$ (6)
+where $\eta=0$ ($\eta=1$) if sites $i$ and $j$ are in the same (different) sublattice, i.e., A or B.
 II.2 Mean-field method in momentum space
-We employ a variational mean-field method in momentum space to contrast the results obtained via the ED method, which has been reported in studying the extended Haldane-Hubbard model without lattice potentials Shao et al.
+A variational mean-field method is employed to analyze the ground-state phase diagram of model (1), which has been reported in studying the extended Haldane-Hubbard model without lattice potentials Shao et al.
 (2021).
-In our case, by introducing the operators $a_{\textbf{k},\sigma}^{{\dagger}}=\frac{1}{\sqrt{N}}\sum_{i\in\text{A}}c_{i,\sigma}^{{\dagger}}e^{\mathrm{i}\textbf{k}\cdot\textbf{r}_{i}}$ and $b_{\textbf{k},\sigma}^{{\dagger}}=\frac{1}{\sqrt{N}}\sum_{i\in\text{B}}c_{i,\sigma}^{{\dagger}}e^{\mathrm{i}\textbf{k}\cdot\textbf{r}_{i}}$, the Hamiltonian (1) can be expressed as
+For our model, by introducing the operators $a_{\textbf{k},\sigma}^{{\dagger}}=\frac{1}{\sqrt{N}}\sum_{i\in\text{A}}c_{i,\sigma}^{{\dagger}}e^{\mathrm{i}\textbf{k}\cdot\textbf{r}_{i}}$ and $b_{\textbf{k},\sigma}^{{\dagger}}=\frac{1}{\sqrt{N}}\sum_{i\in\text{B}}c_{i,\sigma}^{{\dagger}}e^{\mathrm{i}\textbf{k}\cdot\textbf{r}_{i}}$, the Hamiltonian can be expressed as
 $\hat{H}=\hat{H}_{\text{0}}+\hat{H}_{\text{I}},$ (7)
@@
 $\hat{H}_{\rm MF}=\hat{H}_{0}+\sum_{\textbf{k}}\psi_{\textbf{k}}^{{\dagger}}\left(\begin{array}[]{cccc}\varepsilon_{\uparrow}^{a}&\xi_{\uparrow\uparrow}(\textbf{k})&\varepsilon_{\uparrow\downarrow}^{a}&\xi_{\uparrow\downarrow}(\textbf{k})\\ \xi^{*}_{\uparrow\uparrow}(\textbf{k})&\varepsilon_{\uparrow}^{b}&\xi^{*}_{\downarrow\uparrow}(\textbf{k})&\varepsilon_{\uparrow\downarrow}^{b}\\ (\varepsilon_{\uparrow\downarrow}^{a})^{*}&\xi_{\downarrow\uparrow}(\textbf{k})&\varepsilon_{\downarrow}^{a}&\xi_{\downarrow\downarrow}(\textbf{k})\\ \xi^{*}_{\uparrow\downarrow}(\textbf{k})&(\varepsilon_{\uparrow\downarrow}^{b})^{*}&\xi^{*}_{\downarrow\downarrow}(\textbf{k})&\varepsilon_{\downarrow}^{b}\end{array}\right)\psi_{\textbf{k}},$
-where $\psi^{\dagger}_{\textbf{k}}=[a^{\dagger}_{\textbf{k},\uparrow},b^{\dagger}_{\textbf{k},\uparrow},a^{\dagger}_{\textbf{k},\downarrow},b^{\dagger}_{\textbf{k},\downarrow}]$ is the basis for each lattice momentum ${\bf k}$ and
+where $\psi^{\dagger}_{\textbf{k}}=[a^{\dagger}_{\textbf{k},\uparrow},b^{\dagger}_{\textbf{k},\uparrow},a^{\dagger}_{\textbf{k},\downarrow},b^{\dagger}_{\textbf{k},\downarrow}]$ represents the basis for each lattice momentum ${\bf k}$ and
 $\xi_{\sigma\sigma^{\prime}}(\textbf{k})$ $=-\frac{V}{N}\sum_{\textbf{q}}g(\textbf{k}-\textbf{q})\langle b_{\textbf{q},\sigma^{\prime}}^{{\dagger}}a_{\textbf{q},\sigma}\rangle_{\rm MF},$
@@
 The above mean-field equations can be solved self-consistently by making use of the variational mean-field approach.
-Once the free energy has converged, the SDW and CDW order parameters can be obtained by
-$\mathcal{O}_{\text{SDW}}$ $=$ $\left|\frac{1}{2}\left(\langle\vec{S}_{A}\rangle_{\rm MF}-\langle\vec{S}_{B}\rangle_{\rm MF}\right)\right|,$
-$\mathcal{O}_{\text{CDW}}$ $=$ $\left|\left(n^{A}_{\uparrow}+n^{A}_{\downarrow})-(n^{B}_{\uparrow}+n^{B}_{\downarrow}\right)\right|.$ (12)
-Here $\vec{S}_{i}=\frac{1}{2}\sum_{\alpha\beta}c_{i\alpha}^{\dagger}\vec{\sigma}_{\alpha\beta}c_{i\beta}$, and $\vec{\sigma}=(\sigma^{x},\sigma^{y},\sigma^{z})$ is the vector of spin-1/2 Pauli matrices.
+Once the free energy has converged, the SSO and SCO order parameters can be obtained by
+$\mathcal{O}_{\rm SSO}$ $=$ $\left|\frac{1}{2}\left(\langle\vec{S}_{A}\rangle_{\rm MF}-\langle\vec{S}_{B}\rangle_{\rm MF}\right)\right|,$
+$\mathcal{O}_{\rm SCO}$ $=$ $\left|\left(n^{A}_{\uparrow}+n^{A}_{\downarrow})-(n^{B}_{\uparrow}+n^{B}_{\downarrow}\right)\right|.$ (12)
+Here $\vec{S}_{i}=\frac{1}{2}\sum_{\alpha\beta}c_{i\alpha}^{\dagger}\vec{\sigma}_{\alpha\beta}c^{\phantom{\dagger}}_{i\beta}$, and $\vec{\sigma}=(\sigma^{x},\sigma^{y},\sigma^{z})$ is the vector of spin-1/2 Pauli matrices.
 Meanwhile, we use the discrete formulation in its multiband (non-Abelian) version to compute the Chern number Fukui et al.
 (2005).
+Besides, similar to the definitions in Ref. He et al. (2024); Silva et al. , we define the effective potential differences for spin-$\uparrow$ and spin-$\downarrow$ electrons in the MF method as
+$\Delta_{\rm MF}^{\uparrow}$ $=$ $\frac{|\varepsilon_{\uparrow}^{a}-\varepsilon_{\uparrow}^{b}+2\Delta|}{2},$
+$\Delta_{\rm MF}^{\downarrow}$ $=$ $\frac{|\varepsilon_{\downarrow}^{a}-\varepsilon_{\downarrow}^{b}-2\Delta|}{2}.$ (13)
+Notice that $+2\Delta$ in $\Delta_{\rm MF}^{\uparrow}$ and $-2\Delta$ in $\Delta_{\rm MF}^{\downarrow}$ come from the opposite potential differences for spin $\uparrow$ and $\downarrow$ species.
 III Results and analysis
-III.1 Results of the Exact Diagonalization
-In our ED calculations, we choose the $12$A cluster whose reciprocal lattice contains the $\Gamma$, K, ${}^{\prime}$ points, and one pair of M points, see more details in Ref.
+III.1 Results of the exact diagonalization method
+In the ED calculations, we choose the $12$A cluster whose reciprocal lattice encompasses the $\Gamma$, K, K points, and one pair of M points, see more details in Ref.
 Shao et al.
 (2021).
-It has been shown that containing the high-symmetry K and ${}^{\prime}$ points are critical to study the quantum phase transition in interacting Haldane systems Shao et al.
+It has been shown that including the high-symmetry K and K points is crucial for studying the quantum phase transition in interacting Haldane models Shao et al.
 (2021); Varney et al.
 (2010, 2011).
-Meanwhile, we employ the periodic boundary condition and make use of translational symmetries to reduce the Hilbert space size.
-Therefore, discussions with regard to the excited states below are restricted to the momentum subspace where the ground state belongs to.
+The periodic boundary condition is utilized for the calculation of all properties and when determining the Chern numbers, we make use of translational symmetries to reduce the size of the Hilbert space.
 The ($V$, $\Delta$) phase diagrams are shown in Figs.
 1(a), 1(b), 1(c) and 1(d) with $U=0.0$, $U=1.0$, $U=2.0$ and $U=3.0$, respectively.
-The phase transition points (black circles) in Fig.
-1 are identified by the positions of the peaks of fidelity metric $g$.
-Notice that in Fig.
-1(a) with $U=0.0$, the results of Chern number matches the phase diagram obtained from the fidelity metric very well, where we use the blue, green and red squares to represent the Chern number $C=2$, $C=1$ and $C=0$, respectively.
+In Fig.
+1, the phase transition points (black circles) are identified based on the positions of the peaks of fidelity metric $g$.
+Let us first focus on the $U=0.0$ case in Fig.
+1(a), where the Chern number results match the phase diagram obtained from the fidelity metric very well.
+Here we use the blue, green and red squares to represent the Chern number $C=2$, $C=1$ and $C=0$, respectively.
 We can observe from Fig.
-1 (a) that the Chern insulator (CI) phase with $C=2$ is in a region with small $V$ and $\Delta$, the spontaneous symmetry breaking SDW or CDW phases with $C=0$ dominates the system when $\Delta$ or $V$ is large enough.
-Interestingly, an intermediate phase with $C=1$ can be observed surrounded by three expected phases.
-It is different from the phase diagram in Ref.
+1 (a) that the Chern insulator phase with $C=2$ is present in a region with small $V$ and $\Delta$, while the SSO (SCO) phase with $C=0$ dominates the system when $\Delta$ ($V$) is sufficiently large.
+Interestingly, an intermediate phase with $C=1$ can be observed surrounded by the three expected phases.
+This is different from the phase diagram in Ref.
 Vanhala et al.
 (2016), where the $C=1$ phase is sandwiched between a band insulator and Mott insulator.
-Notice that their charge ordered phase (band insulator) is governed by the ionic potential $\Delta_{\rm AB}$ which is spin-independent, and their spin ordered phase (Mott insulator) is governed by the on-site Hubbard interaction $U$.
-While in our case, the charge ordered phase (CDW) is governed by the nearest-neighbor interaction $V$ and the spin ordered phase (SDW) is governed by the spin-dependent lattice potential $\Delta_{\sigma,\alpha}$.
-We speculate the different competition fashions result in the different locations of $C=1$ phase.
-In addition, we observe that the $C=1$ phase in Fig.
-1(a) can exist with very small $V$, while in small $\Delta$ region ($\Delta\leq 0.5$) it vanishes.
-This is completely opposite to our results of MF method in Sec.
-III.2, where the $C=1$ phase can exist in small $\Delta$ region rather than small $V$ region, see Fig.
-3(a).
-We suggest that this is an open issue and expect further studies on it.
-Figure 1: Phase diagram in the parametric space ($V$, $\Delta$) of the model (1) based on the results of fidelity metric $g$ with (a) $U=0.0$, (b) $U=1.0$, (c) $U=2.0$ and (d) $U=4.0$.
+Notice that their charge ordered phase (band insulator) is governed by the ionic potential $\Delta_{\rm AB}$ which is not spin-dependent, and their spin ordered phase (Mott insulator) is governed by the on-site Hubbard interaction $U$.
+In contrast, in our case, the staggered charge order (SCO) phase is governed by the nearest-neighbor interaction $V$ and the staggered spin order (SCO) phase is governed by the spin-dependent lattice potential $\Delta_{\sigma,\alpha}$.
+We propose that the different competition mechanisms result in the different locations of the $C=1$ phase.
+Figure 1: Phase diagram in the parametric space ($V$, $\Delta$) of the model (1) based on the results of fidelity metric $g$ with (a) $U=0.0$, (b) $U=1.0$, (c) $U=2.0$ and (d) $U=3.0$.
 The blue, green and red squares indicate results of Chern number $C=2$, $C=1$ and $C=0$, respectively.
 The black dashed lines in (a) and (d) denote the parameters we choose to show more details below.
-When $U=1.0$ in Fig.
-1(b), however, results of Chern number do not match the phase diagram obtained from the fidelity metric.
-Similar discrepancies between Chern number and fidelity metric have also been discussed in Ref.
-Yuan et al.
-(2023) and they found that increasing $(N_{x},N_{y})$ from (6, 6) to (20, 20) can effectively reduce such discrepancies.
-Remember that $N_{x}$ and $N_{y}$ come from the discretized method to calculate Chern number, see the relevant discussions in Sec.
-II.
-While here we already use $(N_{x}=20,N_{y}=20)$ to calculate the Chern number, and results of $C=1$ (green points) can not even be observed in Figs.
+The impact of the on-site interaction $U$ on the phase diagram is depicted in Figs.
+1(b), 1(c) and 1(d).
+In Fig.
+1(b) with $U=1.0$, there is a discrepancy between the results of the Chern number and the phase diagram obtained from the fidelity metric.
+Moreover, results of $C=1$ (green points) can not even be observed in Figs.
 1(c) and 1(d) with $U=2.0$ and $U=3.0$, respectively.
-We further use $(N_{x}=60,N_{y}=60)$ to check several points in Fig.
-1(d) but with no finding of $C=1$.
-On the other hand, the area of the intermediate phase identified by the fidelity metric decreases but does not vanishes as $U$ increases from $0.0$ to $3.0$.
-(We must have a statement here that in our ED discussions, the $C=1$ phase is equivalent to the intermediate phase when $U=0.0$, and they are not equivalent when $U$ is finite).
+On the other hand, the area of the “intermediate phase” identified by the fidelity metric decreases but does not vanishes as $U$ increases from $0.0$ to $3.0$.
+We discuss the ($U$, $\Delta$) phase diagram with $V=0$ in Appendix A to better understand the influence of $U$ on the ($V$, $\Delta$) phase diagrams.
+We must have a statement here that in our ED discussions, the $C=1$ phase is equivalent to the “intermediate phase” only when $U=0.0$, and they are not equivalent when $U$ is finite because of the discrepancies.
+Similar discussions have also been reported in Refs.
+Shao et al.
+(2021); Yuan et al.
+(2023) and they found that it is better to use the fidelity results to characterize the phase transitions in the interacting Haldane models because the discretized method to calculate Chern number suffers from the finite-size effect more severely.
 The MF results in Sec.
-III.2 also show the existence of $C=1$ in the case of $U=3.0$.
-We thus suggest the issue of whether the $C=1$ phase can exist with $U$ increased to 3.0 may still be open.
-This issue is worth of further study by other methods like Monte Carlo and density matrix renormalization group because in materials $U$ is generally larger than $V$.
-Back to Fig.
-1(a), the situation with $U=0.0$ and finite $V$ may still be realized in ultracold atom systems due to their advances in parameter simulations.
-Figure 2: (a)(e) Four lowest-lying energy levels $E_{\alpha}$, (b)(f) the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$, (c)(g) the structure factors $S_{\text{SDW/CDW}}$, and (d)(h) the fidelity metric $g$ of the model (1) with $V=0.4$ on the left panels and $\Delta=1.4$ on the right panels.
-The on-site interaction $U=0$ and the parameters are corresponding to the balck dashed lines in Fig.
+III.2 also support the existence of $C=1$ in the case of $U=3.0$.
+Even so, we still suggest that the question of whether the $C=1$ phase can exist with $U$ increased to $3.0$ may be open.
+This issue is worth further investigation, especially considering that in materials $U$ is generally larger than $V$.
+Figure 2: (a)(e) Four lowest-lying energy levels $E_{\alpha}$, (b)(f) the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$, (c)(g) the structure factors $S_{\text{SSO/SCO}}$, and (d)(h) the fidelity metric $g$ of the model (1) with $V=0.4$ on the left panels and $\Delta=1.4$ on the right panels.
+The on-site interaction $U=0$ and the parameters are corresponding to the black dashed lines in Fig.
 1(a).
-To detail the critical behaviors of the phase diagram, we select two balck dashed lines with $V=0.4$ and $\Delta=1.4$ in Fig.
-1(a) to calculate the relevant properties.
+To elucidate the critical behaviors of the phase diagram in Fig.
+1(a) with $U=0.0$, we focus on the two black dashed lines with $V=0.4$ and $\Delta=1.4$ to calculate the relevant properties.
+That is, the left and right panels in Fig.
+2 are corresponding to the results of $V=0.4$ and $\Delta=1.4$, respectively.
 The first four lowest-lying energy levels $E_{\alpha}$ ($\alpha=0,1,2,3$ and $\alpha=0$ represents the ground state) are obtained by employing the Arnoldi Lehoucq et al.
@@
 In Fig.
-2(a) with $V=0.4$, the system crosses the $C=2$, $C=1$ and SDW phases as $\Delta$ increases from $0.6$ to $1.4$, and two level crossings between the ground state and one excited state can be observed at $\Delta\approx 0.9$ and $\Delta\approx 1.2$.
+2(a), the system traverses the $C=2$, $C=1$ and SSO phases as $\Delta$ increases from $0.6$ to $1.4$, and two level crossings between the ground state and one excited state are observed at $\Delta\approx 0.9$ and $\Delta\approx 1.2$.
 In Fig.
-2(e) with $\Delta=1.4$, the system crosses the SDW, $C=1$ and CDW phases as $V$ increases from $0.0$ to $2.0$, and two level crossings locates at $V\approx 0.75$ and $V\approx 1.25$.
-It may be more intuitive to exhibit the excitation gaps $\Delta_{\rm ex}^{1}=E_{1}-E_{0}$ and $\Delta_{\rm ex}^{2}=E_{2}-E_{0}$, see Figs.
+2(e), the system crosses the SSO, $C=1$ and SCO phases as $V$ increases from $0.0$ to $2.0$, and two level crossings occur at $V\approx 0.75$ and $V\approx 1.25$.
+To provide a more intuitive representation, we exhibit the excitation gaps $\Delta_{\rm ex}^{(1)}=E_{1}-E_{0}$ and $\Delta_{\rm ex}^{(2)}=E_{2}-E_{0}$, as shown in Figs.
 2(b) and 2(f).
-We can find that the excitation gaps exhibit local minimum values at the critical points where level crossings occur.
-It should be pointed out that in Fig.
-2(f), $\Delta_{\rm ex}^{1}$ becomes very small when $V>1.25$, which is due to the almost degenerate ground state and first excited state in CDW phase.
-Another feature to characterize the phase transitions is the change of the CDW or SDW structure factor ($S_{\text{CDW}}$ or $S_{\text{SDW}}$) in Figs.
+It is evident that the excitation gaps exhibit local minimum values at the critical points where level crossings occur.
+It is worth noting that in Fig.
+2(f), $\Delta_{\rm ex}^{(1)}$ becomes very small when $V>1.25$, owing to the nearly degenerate ground state and first excited state in SCO phase.
+Another feature to characterize the phase transitions is the change of the SCO or SSO structure factor ($S_{\text{SCO}}$ or $S_{\text{SSO}}$), see Figs.
 2(c) and 2(g).
-Values of $S_{\text{CDW}}$ ($S_{\text{SDW}}$) in $C=1$ phase is larger than those in $C=2$ phase but smaller than those in CDW (SDW) phase, which indicating that $C=1$ is an intermediate state as a result of the interplay between topology, electronic correlations and lattice potentials.
-In addition, sharp peaks of the ground-state fidelity metrics $g$ in Figs.
-2(d) and 2(h) can be also observed and used to characterize the critical points.
-Finally, notice that the left and right panels in Fig.
-2 are corresponding to the results of $V=0.4$ and $\Delta=1.4$, respectively.
-For $U=3.0$ in Fig. 1(d), similar calculations and discussions along the black dashed lines of $\Delta=0.4$ and $\Delta=0.6$ are detailed in Appendix A.
+Values of $S_{\text{SCO}}$ ($S_{\text{SSO}}$) in $C=1$ phase is larger than those in $C=2$ phase but smaller than those in the SCO (SSO) phase, indicating again that $C=1$ is an intermediate state resulting from the interplay between topology, electronic correlations and lattice potentials.
+Finally, sharp peaks of the ground-state fidelity metrics $g$ in Figs.
+2(d) and 2(h) can be observed and used to characterize the critical points.
+Besides, for $U=3.0$ in Fig.
+1(d), similar calculations and discussions along the black dashed lines with $\Delta=0.4$ and $\Delta=0.6$ are detailed in Appendix B.
+Before closing this section, we would like to note that the finite-size effect in ED method is discussed in Appendix C.
 III.2 Results of the mean-field method
-To contrast the exact results on small lattice above, we now report the outcomes of the MF method.
-By implementing a variational mean-field approximation, the lattice size can be enlarged significantly.
-We use the $180\times 180$ lattice to calculate the band gap and the order parameters ($\mathcal{O}_{\text{CDW}}$ and $\mathcal{O}_{\text{SDW}}$), while a $30\times 30$ lattice is adopted for the calculations of Chern number.
-Even though their lattice size are different, critical points identified by the Chern number and other properties are consistent, as will be discussed in Fig.
+To contrast the exact results on small lattice above, we now report the outcomes of the mean-field method.
+We use the $180\times 180$ lattice to calculate the band gap and the order parameters ($\mathcal{O}_{\text{SCO}}$ and $\mathcal{O}_{\text{SSO}}$), while a $20\times 20$ lattice is adopted for the calculations of Chern number.
+Despite the differences in lattice size, the critical points identified by the Chern number and other properties are consistent, as will be shown in Fig.
 4.
@@
 3(a), 3(b), 3(c) and 3(d) with $U=0.0$, $U=1.0$, $U=2.0$ and $U=3.0$, respectively.
-Compared to the phase diagrams of ED in Fig.
-1, similar features can be observed: SDW and CDW dominate the system for large $\Delta$ and $V$, respectively, leaving the $C=2$ phase in small ($V$, $\Delta$) region; the $C=1$ phase are surrounded by the above three phases.
-Differently, the increasing interaction $U$ suppresses but does not eliminate the $C=1$ state.
-Figure 3: Mean-field phase diagram in the parametric space ($V$, $\Delta$) of the model (1) based on the results of Chern number $C$ and the order parameters ($\mathcal{O}_{\text{CDW}}$ and $\mathcal{O}_{\text{SDW}}$), with (a) $U=0.0$, (b) $U=1.0$, (c) $U=2.0$ and (d) $U=4.0$.
+In comparison to the phase diagrams of ED in Fig.
+1, similar features can be observed: SSO and SCO dominate the system for large $\Delta$ and $V$, respectively, leaving the $C=2$ phase in small ($V$, $\Delta$) region; the $C=1$ phase are surrounded by the above three phases.
+While differences between the MF and ED phase diagrams include:
+(1) In the MF method, the increased interaction $U$ does not eliminate the $C=1$ phase at least for $U\leq 3.0$.
+(2) The $C=1$ phase in Fig.
+1(a) can exist with very small $V$, while in small $\Delta$ region ($\Delta\leq 0.5$) it vanishes.
+This is opposite to the MF results in Fig.
+3(a), where the $C=1$ phase can exist in small $\Delta$ region rather than small $V$ region.
+Figure 3: Mean-field phase diagram in the parametric space ($V$, $\Delta$) of the model (1) based on the results of Chern number $C$ and the order parameters ($\mathcal{O}_{\text{SCO}}$ and $\mathcal{O}_{\text{SSO}}$), with (a) $U=0.0$, (b) $U=1.0$, (c) $U=2.0$ and (d) $U=3.0$.
 Taking the $U=0.0$ case in Fig.
 3(a) as an example, we show in Figs.
-4(a) and 4(b) the CDW and SDW order parameters ($\mathcal{O}_{\text{CDW}}$ and $\mathcal{O}_{\text{SDW}}$), respectively, as a function of $V$ and $\Delta$.
-We find that $\mathcal{O}_{\text{CDW}}$ vanishes in SDW and $C=2$ phases, while the $C=1$ phase can be regarded as a zone of transition for $\mathcal{O}_{\text{CDW}}$ decreases from $2$ to 00.
-Notice that the phase boundary between SDW and CDW can only be identified from the order parameters, and other phase boundaries are characterized by the results of Chern number, see Fig.
+4(a), 4(b), 4(c) and 4(d) the results of the SCO order parameter ($\mathcal{O}_{\text{SCO}}$), the SSO order parameter ($\mathcal{O}_{\text{SSO}}$), the band gap ($\Delta(\textbf{k})$) and the Chern number ($C$), respectively, as a function of $V$ and $\Delta$.
+It is noteworthy that the phase boundary between SSO and SCO can only be identified from the order parameters, and other phase boundaries are characterized by the results of Chern number, as seen in Fig.
 4(d).
-Even though there are some defect points in Fig.
-4(d) which may be due to the severe quantum fluctuation in the vicinity of the critical points, the phase boundaries can be readily identified.
-For $\mathcal{O}_{\text{SDW}}$ in Fig.
-4(b), the $C=1$ phase is still a transition region with finite values but smaller that those in SDW phase.
+Despite some defect points in Fig.
+4(d), possibly attributed to the severe quantum fluctuation near the critical points, the phase boundaries can be readily identified.
+In Fig.
+4(a), we observe that $\mathcal{O}_{\text{SCO}}$ vanishes in the SSO and $C=2$ phases, while the $C=1$ phase can be regarded as a zone of transition with $\mathcal{O}_{\text{SCO}}$ changing from $0.0$ to $2.0$.
+For $\mathcal{O}_{\text{SSO}}$ in Fig.
+4(b), the $C=1$ phase remains a transition region with finite values but smaller that those in SSO phase.
 These features are similar to the structure factors of ED results shown in Figs.
@@
 4(c).
-The rapid drop or closing of the gap size can be observed at the critical values except for the phase boundary between SDW and CDW, where only discontinuity of gap size appears instead.
+The rapid drop or closing of the gap size can be observed at the critical values, except for the phase boundary between SSO and SCO, where only discontinuity in gap size appears.
 This can be connected with the general picture that the change of a topological invariant is always accompanied by a single-particle gap closing.
-For the phase transition from SDW to CDW side, the Chern number does not change ($C=0$).
-That may be a reason to explain the discontinuity of gap size instead of gap closing, and remember that they are all gapped phases with different gap sizes.
-Figure 4: Contour plots of (a) the CDW order parameter, (b) the SDW order parameter, (c) the band gap and (d) the Chern number, as a function of $V$ and $\Delta$.
+While for the phase transition from SSO to SCO side, the Chern number does not change ($C=0$).
+This may explain the discontinuity of gap size instead of gap closing, considering that the SSO and SCO states are all gapped phases with different gap sizes.
+Figure 4: Contour plots of (a) the SCO order parameter, (b) the SSO order parameter, (c) the band gap and (d) the Chern number, as a function of $V$ and $\Delta$.
 The mean-field approach is adopted for the calculations of the model (1) with $U=0.0$.
+To analyze the origin of the $C=1$ phase, we define the effective potential differences $\Delta_{\rm MF}^{\uparrow}$ and $\Delta_{\rm MF}^{\downarrow}$ in Eq.
+(13), and here we show $\Delta_{\rm MF}^{\uparrow}/t_{2}$ and $\Delta_{\rm MF}^{\downarrow}/t_{2}$ as a function of $V$ in Fig.
+5(a).
+We set $U=0$, $\Delta=0.6$ and $V\in[0.0,1.5]$, see the black dashed line in Fig.
+3(a).
+It is known that in the noninteracting spinless Haldane model, the Chern number $C$ equals $1$ under the condition $|\Delta|/t_{2}<3\sqrt{3}$, and $C$ equals 00 when $|\Delta|/t_{2}>3\sqrt{3}$.
+From Fig.
+5(a) we can observe that:
+$\Delta_{\rm MF}^{\uparrow}/t_{2}<3\sqrt{3}$ and $\Delta_{\rm MF}^{\downarrow}/t_{2}<3\sqrt{3}$ when $C=2$;
+$\Delta_{\rm MF}^{\uparrow}/t_{2}>3\sqrt{3}$ and $\Delta_{\rm MF}^{\downarrow}/t_{2}<3\sqrt{3}$ when $C=1$;
+$\Delta_{\rm MF}^{\uparrow}/t_{2}>3\sqrt{3}$ and $\Delta_{\rm MF}^{\downarrow}/t_{2}>3\sqrt{3}$ when $C=0$.
+Here the vertical dashed lines in Fig.
+5(a) separate the $C=2$, $C=1$ and $C=0$ regions.
+In Fig.
+5(b), we show the band gap $\Delta(\textbf{k})$ as a function of $V$, where two gap closures can be found at the critical points, denoted by the vertical dashed lines identical to those in Fig.
+5(a).
+These features indicates that the formation of the $C=1$ phase is due to the spontaneous SU(2) symmetry breaking, with only one spin species in topological state.
+Figure 5: (a) The effective potential differences in units of $t_{2}$ $\Delta_{\rm MF}^{\uparrow}/t_{2}$ and $\Delta_{\rm MF}^{\downarrow}/t_{2}$, and (b) the band gap as a function of $V\in[0.0,1.5]$ (see the black dashed line in Fig.
+3(a)).
+Other parameters: $U=0$ and $\Delta=0.6$.
+The horizontal dashed line in (a) denotes the value of $3\sqrt{3}$ in y axis, and the vertical dashed lines in (a) and (b) separate the $C=2$, $C=1$ and $C=0$ regions.
 IV Summary and Discussion
-To summarize, we studied the interacting spinful Haldane model at half-filling on the honeycomb lattice with spin-dependent sublattice potentials.
-By employing the exact-diagonalization (ED) and mean-field (MF) method, we obtained the similar ground-state phase diagrams with regard to the nearest-neighbor interaction $V\in[0,3]$ and sublattice potential difference $\Delta\in[0,3]$: the CDW and SDW phases are in the large $V$ and $\Delta$ region, respectively; the $C=2$ phase dominates the system when both $V$ and $\Delta$ are small enough; a intermediate phase with $C=1$ are surrounded by the above three phases.
-Except for the change of topological invariant, i.e., the Chern number, other features like the closure of excitation gap, change of structure factors and peak of the fidelity metric in ED results are also observed.
-As a contrast, the mean-field Chern number, band gap and local order parameters are also studied, which makes the existence of such $C=1$ phase more evident.
-On the other hand, the interaction $U$ suppresses the $C=1$ phase but leaving an open question of whether it can exist with $U$ increased to $3.0$.
-In ED, $C=1$ can not be found when $U=3.0$ even though an intermediate phase can still be identified by results of the fidelity metric.
-In MF, $C=1$ phase preserves for $U=3.0$ even though its area is very small.
-Further studies on this interesting issue are thus expected.
+In summary, we investigated the interacting spinful Haldane model at half-filling on the honeycomb lattice with spin-dependent sublattice potentials.
+By employing the exact-diagonalization (ED) and mean-field (MF) methods, we obtained similar ground-state phase diagrams for the nearest-neighbor interaction $V\in[0,3]$ and sublattice potential difference $\Delta\in[0,3]$.
+The staggered spin order (SSO) and staggered charge order (SCO) phases dominate in the large $\Delta$ and $V$ region, respectively; the $C=2$ phase prevails when both $V$ and $\Delta$ are small enough; and an intermediate phase with $C=1$ is surrounded by the aforementioned three phases.
+Except for the change in the topological invariant, other features such as the closure of the excitation gap, changes in structure factors, and peaks in the fidelity metric were also observed in the ED results.
+Meanwhile, we examined the mean-field Chern number, band gap and local order parameters, also providing clear evidence for the existence of the $C=1$ phase.
+Especially, by analyzing the effective potential differences $\Delta_{\rm MF}^{\uparrow}$ and $\Delta_{\rm MF}^{\downarrow}$, the origin of this exotic phase is also attributed to the spontaneous SU(2) symmetry breaking.
+However, the interaction $U$ suppresses the $C=1$ phase, raising an open question about its potential existence as $U$ increased to $3.0$.
+In the ED approach, $C=1$ can not be found when $U=3.0$ even though the “intermediate phase” can still be identified from the fidelity metric results.
+In the MF approach, the $C=1$ phase preserves for $U=3.0$, albeit with a very small area.
+Finally, we would like to note that in previous investigations, the staggered spin order (SSO) has been denoted as the antiferromagnetic (AF)Jiang et al.
+(2018); Wang and Qi (2019) or spin-density-wave (SDW) state He et al.
+(2011, 2024); Silva et al.
+; Yuan et al.
+(2023); Shao et al.
+(2021), and the staggered charge order (SCO) has been labeled as the band insulator (BI)Wang and Qi (2019); Vanhala et al.
+(2016) or charge-density-wave (CDW)Wang and Qi (2019); He et al.
+(2024); Silva et al.
+; Yuan et al.
+(2023); Shao et al.
+(2021) state.
+While in this paper, we refer to them as SSO and SCO to stress their staggered characteristics on A and B sublattices.
 Acknowledgements.
+The authors acknowledge insightful discussions with E.
+V.
+Castro, R.
+Mondaini, H.
+Lu and S.
+Hu.
 C.
@@
 2022YFA1402704).
-Appendix A details of some properties with $U=3$
-Figure 5: (a)(e) Four lowest-lying energy levels $E_{\alpha}$, (b)(f) the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$, (c)(g) the structure factors $S_{\text{SDW/CDW}}$, and (d)(h) the fidelity metric $g$ of the model (1) with $\Delta=0.4$ on the left panels and $\Delta=0.6$ on the right panels.
-The on-site interaction $U=3$ and the parameters are corresponding to the balck dashed lines in Fig.
+Appendix A the ED phase diagram in the parametric space ($U$, $\Delta$) with $V=0$
+Figure 6: (a) The exact-diagonalization (ED) phase diagram of the model (1) in the parametric space($U$, $\Delta$) with $V=0$.
+The black dashed line denotes the parameters we choose to show the structure factor $S_{\text{SSO}}$ in (b), the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$ in (c), and the fidelity metric $g$ in (d).
+To better understand the impact of the interaction parameter $U$ on the phase diagram, we present the ED phase diagram of our model as a function of $U$ and $\Delta$ in Fig.
+6(a), with $V=0$.
+It can be observed that both $U$ and $\Delta$ favor the SSO order, and $\Delta$ is more efficient in suppressing the $C=2$ phase compared to $U$.
+Specifically, for $\Delta=0$, the critical point is located at $U=8.1$, while for $U=0$, the critical point is located at $\Delta=1.03$.
+This explains why increasing $U$ from 00 to $3$ does not eliminate the intermediate phase in the ($V$, $\Delta$) phase diagram, as depicted in Fig.
+1 of the main text.
+Furthermore, we choose $U=4$ [black dashed line in Fig.
+6(a)] to plot the structure factor of SSO in Fig.
+6(b), the excitation gaps in Fig.
+6(c) and the fidelity metric in Fig.
+6(d).
+The rapid change of $S_{\text{SSO}}$, the closure of $\Delta_{\text{ex}}^{(1)}$ and the sharp peak of the fidelity metric $g$ can be observed, respectively.
+Notably, no intermediate phase can be identified from these properties, indicating that the $C=1$ phase results from the competition between $V$ and $\Delta$.
+Appendix B details of some properties in ED method with $U=3$
+Figure 7: (a)(e) Four lowest-lying energy levels $E_{\alpha}$, (b)(f) the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$, (c)(g) the structure factors $S_{\text{SSO/SCO}}$, and (d)(h) the fidelity metric $g$ of the model (1) with $\Delta=0.4$ on the left panels and $\Delta=0.6$ on the right panels.
+The on-site interaction $U=3$ and the parameters are corresponding to the black dashed lines in Fig.
 1(d).
-In the case of $U=3.0$, we select $\Delta=0.4$ (left panel) and $\Delta=0.6$ (right panel) to present the relevant properties in Fig.
-5.
-Remember that we do not observe the results of $C=1$ in Fig.
-1(d) with $U=3.0$.
-The line of $\Delta=0.4$ ($\Delta=0.6$) crosses the $C=2$ phase (SDW phase), the intermediate phase and the CDW phase as $V$ increases from $1.0$ to $2.0$ [see black dashed lines in Fig.
+In Fig.
+7, considering the case of $U=3.0$, we choose $\Delta=0.4$ (left panel) and $\Delta=0.6$ (right panel) to illustrate the relevant properties.
+The line of $\Delta=0.4$ ($\Delta=0.6$) in Fig.
+1(d) crosses the $C=2$ phase (SSO phase), the “intermediate phase” and the SCO phase as $V$ increases from $1.0$ to $2.0$ [see black dashed lines in Fig.
 1(d)].
-From the ground-state fidelity metrics $g$ in Figs.
-5(d) and 5(h), we can find the sharp peaks between the intermediate and the CDW phases.
-Accompanied with the shape peaks, level crossings between the ground state and second excited state [Figs.
-5(a) and 5(e)], minimum of the second excitation gap $\Delta_{\text{ex}}^{(2)}$ [Figs.
-5(b) and 5(f)] and sudden changes of structure factors [Figs.
-5(c) and 5(g)] can be observed.
-Notice that in CDW phase the ground state and first excited state are almost degenerate.
-On the other hand, only “hump”s can be observed at the phase boundaries between the $C=2$ (or SDW) and the intermediate phase.
-Near the positions of the “hump”s, evolutions of the low-energy levels are more complicated and changes of structure factors are more smooth.
-Such features have also been discussed in Refs.
+Examining the fidelity metrics $g$ in Figs.
+7(d) and 7(h), we observe sharp peaks between the “intermediate phase” and the SCO phase, located at $V\approx 1.5$ and $V\approx 1.6$, respectively.
+In the vicinity of the locations of the shape peaks, level crossings between the ground state and the second excited state can be observed [Figs.
+7(a) and 7(e)], as well as the minimum values of the second excitation gap $\Delta_{\text{ex}}^{(2)}$ [Figs.
+7(b) and 7(f)] and sudden changes in structure factors [Figs.
+7(c) and 7(g)].
+Notice that in SCO phase the ground state and first excited state are nearly degenerate.
+On the other hand, only “hump”s are observed at the phase boundaries between the $C=2$ (or SSO) and the “intermediate phase”, locating at $V\approx 1.32$ in Fig.
+7(d) [$V\approx 1.27$ in Fig.
+7(h)].
+Near the positions of the “hump”s, level crossings between the ground state and the first excited state [Figs.
+7(a) and 7(e)], the local minimum values of the first excitation gap $\Delta_{\text{ex}}^{(1)}$ [Figs.
+7(b) and 7(f)] and smooth changes in structure factors [Figs.
+7(c) and 7(g)] can be observed.
+Similar features have been discussed in Refs.
 Shao et al.
 (2021); Yuan et al.
-(2023), where its reason is attributed to the finite-size effect and it can be eliminated by using different clusters or twisted boundary conditions.
+(2023), where the reason is attributed to the finite-size effect, and it can be mitigated by employing different clusters or twisted boundary conditions.
+Appendix C the ED phase diagram in the parametric space ($V$, $\Delta$) on a 6-site lattice
+Figure 8: (a) The exact-diagonalization (ED) phase diagram of the model (1) in the parametric space($V$, $\Delta$) with $U=0$, on a $6$-site cluster shown in (c).
+The black dashed line in (a) denotes the parameters we choose to show the spectrum in (b), the excitation gap $\Delta_{\text{ex}}^{(2)}$ in (c), and the structure factors $S_{\rm SSO/SCO}$ in (d).
+In the beginning of Sec.
+III, we mentioned that we utilized a $12$A cluster, whose reciprocal lattice contains the K points Shao et al.
+(2021), to study the phase transitions of the model (1) in the main text.
+Now we would like to discuss the finite-size effect in our ED study.
+Is is known that during the phase transition of noninteracting Haldane model, the gap closes at the K points.
+It’s important to note that the condition for reciprocal lattice of a cluster to contain the K points can only be satisfied when the site number of the cluster is a multiple of $6$Varney et al.
+(2010).
+This condition has been demonstrated to be crucial for characterizing the topological phase transitions in the interacting Haldane modelsVarney et al.
+(2010, 2011); Shao et al.
+(2021); Yuan et al.
+(2023).
+Unfortunately, due to limitations in our current computational resources, we are unable to access an $18$A cluster.
+Therefore, we conducted our phase diagram analysis on the $6$A cluster instead, as shown in subplot of Fig.
+8(c).
+As expected, an intermediate phase can be observed in the ($V$, $\Delta$) phase diagram in Fig.
+8(a) and we select the dashed line with $\Delta=1.5$ to illustrate more detailed properties.
+In Fig.
+8(b), we depict four lowest-lying energy levels, and the level crossings at two critical points can be observed.
+It’s worth noting that the ground state $E_{0}$ and first excited state $E_{1}$ are degenerate in the SCO and intermediate phase, while the first excited state $E_{1}$ and the second excited state $E_{2}$ are degenerate in the SSO phase.
+As a result, it is more appropriate to use the second excitation gap, $\Delta_{\rm ex}^{(2)}$, to characterize the gap closure at phase transition points, as shown in Fig.
+8(c).
+Additionally, we present the structure factors $S_{\rm SSO}$ and $S_{\rm SCO}$ in Fig.
+8(d), where the transition nature of the intermediate phase is clearly visible.